ADNAN MUHAMMED

Grade 12,

A partical thrown up vertically with certain initial velocity, reaches maximum height of 100m.Another particle of double mass is thrown up with double velocity and reaches a maximum height of.

A partical thrown up vertically with certain initial velocity, reaches maximum height of 100m.Another particle of double mass is thrown up with double velocity and reaches a maximum height of.

Grade:12th pass

1 Answers

Susmita
5 years ago
Let 1st particle have an initial velocity u.
At maximum height velocity will be zero.
v=u-gt=0
Or,t=u/g
Height
h=ut-gt2/2=u2/g-gu2/2g2=u2/g-u2/2g=u2/2g=100
In the 2nd case,u'=2u.so 
v'=u'-gt'=0
Or,t'=u'/g
h'=u't-gt'2/2=u'2/2g=4u2/2g=4h=400m